• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nanda

    Log entry time 12:50:05 on March25,2005

    Entry number 142392

    keyword=Compton Rates

    MCC carried out a test with slits fully open to see the effect of compton counting rates. Electron rates jumped from 500 to 6000 Hz/uA. A beam tail was confirmed.

    This tail was arising from injector chopper phase mismatch.
    MCC reworked the injector tune by adjusting chopper phase. After the retune (15 mins) compton rates were back down to acceptable levels (600 Hz) with slits open Fig.1.

    We agreed to run through the weekend, with new injector tune and slits out. Hall A should monitor compton rates. If compton rates degrade beyond a couple kHz/uA, use slits (setting 50) for a quick fix to clean up beam.






    FIGURE 1